# Do Not Worry if You Do Not Have Data: Building Pretrained Language Models Using Translationese

### v2

@@ B.1 Training
B.2 Extended pretraining
For the mini-1k models, we randomly sample 100M tokens from the clean subset of IndicMonoDoc for the target language, and for the base-1k model, we sample 200M for extended pretraining. We use the same hyperparameters as training and perform extended pretraining for 2 epochs over this newly sampled clean data.
B.3 Fine-tuning
For GLUE tasks we use the dev split on the clean part and do hyperparameter tuning to achieve the best scores, and then we use the same hyperparameters for all synthetic experiments. For IndicGLUE we follow a similar setting for the val split to find good hyperparameters and report results on the test split like Kakwani et al. (2020). For all classification and regression tasks, we use a single linear layer and use an appropriate activation function for classification and regression respectively. We use an Adam optimizer Kingma and Ba (2015) with a learning rate of $1e^{-5}$ and a batch size of 48. For NLG tasks we do extended pretraining using a separator token in between the input and output sequence with an effective batch size of 768 examples and only calculate loss for the output sequence. We use an AdamW optimizer Loshchilov and Hutter (2019) with learning rate = $6e^{-4}$, weight decay = $1e^{-1}$, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95 and $\epsilon$ = $1e^{-5}$. For translation, we randomly sample 1M parallel sentence for each language pair from the samanantar corpus Ramesh et al. (2022) and evaluate on FloRes Team et al. (2022), IN22-Conv and IN22-Gen Gala et al. (2023). We list the batch size and number of epochs of each task in Table 11.
@@ B.3 Fine-tuning
B.4 Evaluation
We use torch metrics [^12] https://lightning.ai/docs/torchmetrics/stable/pages/lightning.html to calculate accuracy, f1-score, Pearson correlation, Matthew’s correlation coefficient. We report chrF++ scores131313chrF++ signature
@@ B.4 Evaluation
and BLEU scores141414sacreBLEU signature:
@@
Appendix C Perplexity filtering
@@ C.1 Creating synthetic data
“Translationese” is a term used to describe peculiarities in the text translated into a specific language, differentiating it from content originally written in that language Gellerstam (1986). Translated texts into the target language (via humans or machine-generated) often show distinctive features that differentiate them from their original counterparts in the target language. These disparities arise from either the influence of the translation process itself on the final product or the inherent “fingerprints” of the source language subtly present in the target language rendition Rabinovich and Wintner (2015). This is a common phenomenon in translation models where the target language translations often show characteristics of the source language and add bias to the evaluation of downstream tasks Toral et al. (2018), Zhang and Toral (2019), Graham et al. (2019). So far a lot of work on synthetic translated data has been done for using back translations Sennrich et al. (2016a), Edunov et al. (2018) for improving Machine translation performance Marie et al. (2020),Bogoychev and Sennrich (2019),Ni et al. (2022) or for classification tasks like native language identification Goldin et al. (2018), etc. Tranlationese data has been used for many tasks but we explore the efficacy of using translationese data for pretraining of language models. We collect monolingual corpora in the source language as mentioned in section 4 and utilize a powerful off-the-shelf translation model IndicTrans2 Gala et al. (2023) to generate translationese data. We use the 1B en-indic version [^17] https://huggingface.co/ai4bharat/indictrans2-en-indic-1B of IndicTrans2 using beam search with a beam value of 5 to translate 5 billion English tokens from IndicMonoDoc to Hindi and Gujarati. Since IndicTrans2 can only handle a max sentence length of 256 BPE tokens, we split the documents using Moses Sentence Splitter [^18] https://pypi.org/project/mosestokenizer/ to perform translations into the target language at the sentence level and then merge again to form documents. We also repair translations that exceed in length 256 BPE tokens using the TinyLM trained on clean corpora as mentioned in Section 5 to complete the sentence translation, we encounter only 0.002% of such cases. We also do the reverse using the 1B indic-en version [^19] https://huggingface.co/ai4bharat/indictrans2-indic-en-1B of IndicTrans2 to translate 5B Hindi tokens and 900M Gujarati tokens from IndicMonoDoc to English. Together these make up the unfiltered synthetic translationese data in English, Hindi, and Gujarati. We use this corpus for the synthetic and clean+synthetic part of our experiments.
@@ C.2 Perplexity filtering
Following Figure 2, we use these TinyLMs to filter the generated synthetic translationese corpora from IndicTrans2. We do this by using perplexity as a measure of document quality score. For language models, perplexity quantifies how well a model predicts a sequence of tokens. A lower perplexity indicates better predictive performance. It is calculated by:
@@
where the negative log-likelihood measures the error of the model’s predictions. While calculating perplexity over a sequence of tokens $W\in$ $w_{1},w_{2},\ldots,w_{N}$ we skip the first $s$ tokens where $s=10$, $e=1024$ and calculate loss until only the first $e$ tokens of the document. We find setting $e$ to larger values can lead to higher variance in the document scores due to the size of the TinyLM. After initial analysis, we choose $s$ and $e$ such that we remove the high uncertainty of the language at the start of an unseen document and avoid penalizing longer documents due to the fragility of the extrapolation ability of TinyLM [^20] During experiments we saw that these TinyLMs can only go up to a certain context length before deteriorating in quality.. Note that it is important to choose $e$ such that the language model gives a uniform estimate of perplexity over an already seen sequence of tokens $\in$ $w_{s},w_{s+1},\ldots,w_{e}$. For our experiments, we use the TinyLMs to score all synthetically generated translationese data and calculate a document score using the above method. Following Laurençon et al. (2022) we do subsampling by thresholding document perplexity scores except Laurençon et al. (2022) did it using Ken-LM Heafield (2011) and we do it using our TinyLM. We keep the threshold value such that we include enough documents to reach the computed optimal token count for pretraining experiments.
Language IndicCorpv2 Ours
bn 926.00 5258.47
en 6501.00 11986.53
gu 901.00 887.18
hi 6107.00 11268.33
kn 875.00 567.16
ml 931.00 845.32
mr 795.00 1066.76
ne 852.00 1542.39
pa 732.00 449.61
ta 476.00 2171.92
te 731.00 767.18
ur 667.00 2391.79
as 67.00 57.64
brx 2.50 2.25
doi 0.10 0.37
gom 31.90 2.91
kas 0.06 1.27
mai 13.70 1.51
mni 0.60 0.99
or 122.00 81.96
sa 125.00 80.09
sat 4.00 3.05
sd 13.20 83.81
Table 12: Languagewise corpora size comparison in Million tokens
Appendix D IndicMonoDoc
@@ D.1 Crawling
D.2 Post processing
A lot of crawled content consists of unwanted text like HTML tags, emoticons, and text in another language. We use manual filtering pipelines inspired by OSCAR Ortiz Suárez et al. (2019), Abadji et al. (2022) to remove such content. We additionally use a language detection-based (LID) filtering using cld3232323https://github.com/google/cld3 and IndicLID-FTN model Madhani et al. (2023a) to discard languages not of interest. Following Doddapaneni et al. (2023) we perform document filtering to remove offensive text from the corpora using a list of offensive words and phrases extended from work by Team et al. (2022) which consists of offensive words in 209 languages. We also use a Romanized version of this list using the transliteration tool by Madhani et al. (2023b) to perform toxic document filtering in 17 languages. Following Kakwani et al. (2020) & Doddapaneni et al. (2023) we merge all the filtered corpus with Wikipedia, OSCAR Ortiz Suárez et al. (2019) and some dumps of mC4 Xue et al. (2021). Finally, we perform deduplication at paragraph level using Murmurhash algorithm [^24] https://pypi.org/project/mmh3/ with a 128-bit unsigned hash for each monolingual split of the corpora. After all post-processing steps, the language wise size of the corpora is mentioned in Table 12. A major chunk of the corpus is comprised of English, Hindi, and Bengali which make up 72.15% of the corpora.

### v3

@@ B.2 Training
B.3 Extended pretraining
For the mini-1k models, we randomly sample 100M tokens from the clean subset of IndicMonoDoc for the target language, and for the base-1k model, we sample 200M for extended pretraining. We use the same hyperparameters as training and perform extended pretraining for 2 epochs over this newly sampled clean data. For scaling experiments, we utilize TorchTune [^17] https://github.com/pytorch/torchtune for fine-tuning Llama-3-8B and Gemma-2B models. For a fair comparison, we limit each data split to 344M tokens for Gujarati and 465M tokens for Marathi and follow a similar procedure as described in Section 3.4 to generate and filter data for Marathi. We perform extended training for a single epoch using LoRA (Hu et al., 2021) finetuning on $W_{q}$, $W_{v}$ projection matrices using $\alpha$=$16$ and $r$=$8$. We keep the learning rate at $3e^{-5}$ with a weight decay of $0.01$ and an effective batch size of 58k. We use the AdamW optimizer Loshchilov and Hutter (2019) with 1000 warmup steps and a cosine learning rate scheduler.
B.4 Fine-tuning
For GLUE tasks we use the dev split on the clean part and do hyperparameter tuning to achieve the best scores, and then we use the same hyperparameters for all synthetic experiments. For IndicGLUE we follow a similar setting for the val split to find good hyperparameters and report results on the test split like Kakwani et al. (2020a). For all classification and regression tasks, we use a single linear layer and use an appropriate activation function for classification and regression respectively. We use an Adam optimizer Kingma and Ba (2015) with a learning rate of $1e^{-5}$ and a batch size of 48. For NLG tasks we do extended pretraining using a separator token in between the input and output sequence with an effective batch size of 768 examples and only calculate loss for the output sequence. We use an AdamW optimizer Loshchilov and Hutter (2019) with learning rate = $6e^{-4}$, weight decay = $1e^{-1}$, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95 and $\epsilon$ = $1e^{-5}$. For translation, we randomly sample 1M parallel sentence for each language pair from the samanantar corpus Ramesh et al. (2022) and evaluate on FloRes Team et al. (2022), IN22-Conv and IN22-Gen Gala et al. (2023). We list the batch size and number of epochs of each task in Table 20.
B.5 Prompting
We use random sampling from validation sets whenever available and utilize other examples from the test set otherwise. We take 5 random samples for each evaluation of NLG tasks and 10 random samples for each evaluation of NLU tasks. We list down the prompt used below for Marathi evaluations, we use similar prompts for Gujarati as well.
[…]
@@ B.5 Prompting
B.6 Evaluation
We use torch metrics [^18] https://lightning.ai/docs/torchmetrics/stable/pages/lightning.html to calculate accuracy, f1-score, Pearson correlation, Matthew’s correlation coefficient. We report chrF++ scores191919chrF++ signature
@@ B.6 Evaluation
and BLEU scores202020sacreBLEU signature:
@@
Appendix C Utilising Translationese
In this section we provide details on the generation and filtering of translationese data for our experiments.
@@ C.1 Creating synthetic data
“Translationese” is a term used to describe peculiarities in the text translated into a specific language, differentiating it from content originally written in that language Gellerstam (1986). Translated texts into the target language (via humans or machine-generated) often show distinctive features that differentiate them from their original counterparts in the target language. These disparities arise from either the influence of the translation process itself on the final product or the inherent “fingerprints” of the source language subtly present in the target language rendition Rabinovich and Wintner (2015). This is a common phenomenon in translation models where the target language translations often show characteristics of the source language and add bias to the evaluation of downstream tasks Toral et al. (2018); Zhang and Toral (2019); Graham et al. (2019). So far a lot of work on synthetic translated data has been done for using back translations Sennrich et al. (2016a); Edunov et al. (2018) for improving Machine translation performance Marie et al. (2020); Bogoychev and Sennrich (2019); Ni et al. (2022) or for classification tasks like native language identification Goldin et al. (2018), etc. Tranlationese data has been used for many tasks but we explore the efficacy of using translationese data for pretraining of language models. We collect monolingual corpora in the source language as mentioned in Section 3.1 and utilize a powerful off-the-shelf translation model IndicTrans2 Gala et al. (2023) to generate translationese data. Since IndicTrans2 can only handle a max sentence length of 256 BPE tokens, we split the documents using Moses Sentence Splitter [^23] https://pypi.org/project/mosestokenizer/ to perform translations into the target language at the sentence level and then merge again to form documents. We also repair translations that exceed in length 256 BPE tokens using the TinyLM trained on clean corpora as mentioned in Section 4 to complete the sentence translation, we encounter only 0.002% of such cases. We use this corpus for the synthetic and clean+synthetic part of our experiments.
@@ C.2 Perplexity filtering
Following Figure 2, we use these TinyLMs to filter the generated synthetic translationese corpora from IndicTrans2. We do this by using perplexity as a measure of document quality score. For language models, perplexity quantifies how well a model predicts a sequence of tokens. A lower perplexity indicates better predictive performance. While calculating perplexity over a sequence of tokens $W\in$ $w_{1},w_{2},\ldots,w_{N}$ we skip the first $s$ tokens where $s=10$, $e=1024$ and calculate loss until only the first $e$ tokens of the document. We find setting $e$ to larger values can lead to higher variance in the document scores due to the size of the TinyLM. After initial analysis, we choose $s$ and $e$ such that we remove the high uncertainty of the language at the start of an unseen document and avoid penalizing longer documents due to the fragility of the extrapolation ability of TinyLM [^24] During experiments we saw that these TinyLMs can only go up to a certain context length before deteriorating in quality.. Note that it is important to choose $e$ such that the language model gives a uniform estimate of perplexity over an already seen sequence of tokens $\in$ $w_{s},w_{s+1},\ldots,w_{e}$. For our experiments, we use the TinyLMs to score all synthetically generated translationese data and calculate a document score using the above method. Following Laurençon et al. (2022), we do subsampling by thresholding document perplexity scores except Laurençon et al. (2022) did it using Ken-LM Heafield (2011) and we do it using our TinyLM. We keep the threshold value such that we include enough documents to reach the computed optimal token count for pretraining experiments.
High Resource Low Resource
Lang #Tokens Lang #Tokens
bn 5,258.47 as 57.64
en 11,986.53 brx 2.25
gu 887.18 doi 0.37
hi 11,268.33 gom 2.91
kn 567.16 kas 1.27
ml 845.32 mai 1.51
mr 1,066.76 mni 0.99
ne 1,542.39 or 81.96
pa 449.61 sa 80.09
ta 2,171.92 sat 3.05
te 767.18 sd 83.81
ur 2,391.79
Table 21: Languagewise corpora size in Million tokens
Appendix D Qualitative Analysis
Since translation errors occur frequently, leading to biased, ungrammatical, or erroneous translations that can have drastic consequences on training, it is important to mitigate or remove such errors in translationese corpora. The most common machine translation errors include mistranslations due to ambiguous words, incorrect handling of expressions, syntax and grammar errors, and issues with preserving context across longer sentences. Many approaches have been proposed to address these errors, but most are computationally expensive, especially when translationese data is used for pretraining. Instead, we ask whether a language model can identify such issues. To investigate this, we examine which types of English sentences were filtered.
In many cases, we found that the filtered documents included errors like code-mixing and repetitions, often generated by the diverging output of the translation model. This is expected since such phenomena are rarely seen in natural written language, and the model assigning high entropy suggests an unlikely sequence outcome. Although the model regarded some erroneous instances as false positives, many such cases were successfully avoided. We also noticed that, due to the small size of the model, many perfectly good documents were filtered out because of the model’s inability to evaluate them. This issue was observed less frequently in larger models used for evaluation. As seen in Figure 8, the filtering model fails to understand complex words and named entities, which it regarded as unlikely due to its preference for simpler terms and more likely entities. Other common elements that were filtered included numbers, dates, and abbreviations. While this can lead to the loss of valuable information, as many good documents are discarded, we empirically observe the benefits of such filtering. These errors could likely be reduced by using a larger filtering model that can better approximate the source language but we leave this analysis for future work.
Figure 8: Heatmap of perplexity over filtered sentences.
Appendix E IndicMonoDoc
@@ E.1 Crawling
E.2 Post processing
A lot of crawled content consists of unwanted text like HTML tags, emoticons, and text in another language. We use manual filtering pipelines inspired by OSCAR Ortiz Suárez et al. (2019), Abadji et al. (2022) to remove such content. We additionally use a language detection-based (LID) filtering using cld3272727https://github.com/google/cld3 and IndicLID-FTN model Madhani et al. (2023a) to discard languages not of interest. Following Doddapaneni et al. (2023) we perform document filtering to remove offensive text from the corpora using a list of offensive words and phrases extended from work by Team et al. (2022) which consists of offensive words in 209 languages. We also use a Romanized version of this list using the transliteration tool by Madhani et al. (2023b) to perform toxic document filtering in 17 languages. Following Kakwani et al. (2020a) & Doddapaneni et al. (2023) we merge all the filtered corpus with Wikipedia, OSCAR Ortiz Suárez et al. (2019) and some dumps of mC4 Xue et al. (2021). Finally, we perform deduplication at paragraph level using Murmurhash algorithm [^28] https://pypi.org/project/mmh3/ with a 128-bit unsigned hash for each monolingual split of the corpora. After all post-processing steps, the language wise size of the corpora is mentioned in Table 21. A major chunk of the corpus is comprised of English, Hindi, and Bengali which make up 72.15% of the corpora.